OpenStudy (anonymous):

if a = i + j + 5k and b = i + j + 4k. Compute the cross product a x b.

OpenStudy (anonymous):

i come up with 9i -9j + 2k and that isn't correct...can someone show me where i'm going wrong here?

OpenStudy (anonymous):

did you try this formula? a x b=(a2b3-a3b2)i -(a1b3-a3b1)j +(a1b2-a2b1)k

OpenStudy (anonymous):

did you get the answer sonofa?

OpenStudy (anonymous):

axb=(4-5)i-(4-5)j+(1-1)k axb=-i +j ans

OpenStudy (anonymous):

lol....i fell alseep...sorry